OpenStudy (anonymous):

-2/5v=8 Solve for v and simplify your answer as much as possible.

OpenStudy (anonymous):

v=-1/20

OpenStudy (lgbasallote):

cross multiply -2/5 = 8v now divide both sides by 8... -2/40 = v simplify... -1/20 = v